• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Snyder

    Log entry time 08:00:25 on July27,2005

    Entry number 149634

    keyword=Owl Shift Summary

    00:00   Hachemi and Rob down in hall to reboot ROCs, and VME crate 
            needed to be restarted
    
    00:05   Run 3856 started. Pedestal until approx. 43,000 and then 
            production
    
    01:10  Run 3857  - Q^2 scan preformed by Kirsten F.
    
    01:26  Run 3858 - Production 
    
    02:23  Run 3859 - Production
    
    03:19  Run 3860 - Production
    
    04:15  Run 3861 - Production
    
    05:12  Run 3862 - Production but during most of the run there was no
                      beam
    
    06:09  Run 3863 - Production 
    
    07:06  Run 3864 - Production